1. Express the following signals in their trigonometric, polar and compact exponential forms. if it is necessary to expand the signals and indicate the fundamental frequency \omega_{0} of them.

e) a (t) cos2t f) x (t) = sin? t 9) (t)cos3t h)(t)n3 t
Answer #1

e) x(t) = cos2 t

= 1 + cos 2t / 2

= ;(1 + cos2t)

= 2 (1

f) x(t) = sin2 t

= 1 - cos 2t / 2

= 2 1- cos2t)

=  2 (1

g) x(t) = cos3t

= cos 3t - 3cost / 4

h) x(t) = sin3t

= 3sint -sin 3t / 4